Find the slope of the line that passes through the given pair of points. (If an answer is undefined, enter UNDEFINED.)
(−a + 3, b − 1) and (a + 3, −b)
Please help. Ive tried almosst everything!!

Found 2 solutions by nerdybill, mananth:
Answer by nerdybill(7384)

(−a + 3, b − 1) and (a + 3, −b)
.
slope between two points:
m = (y2 - y1)/(x2 - x1)
plug in the given data:
m = (-b - (b − 1))/((a+3 - (−a + 3))
m = (-b + b + 1)/((a+3 + a - 3)
m = (1)/(2a)
m (slope) = 1/(2a)

Answer by mananth(16946)

(-a+3, b-1) ( a+3,-b)
slope = (y2-y1)/(x2-x1)
m=(-b-(b-1))/((a+3)-(-a+3))
m=(-2b+1)/2a